OpenStudy (anonymous):

Triangle GHI is similar to triangle JKL. If IH= 36, KP=18, and LK= 24, then MH= ____? A. 12 B. 20 C. 27 D. 48

OpenStudy (anonymous):

@_greatmath7

OpenStudy (anonymous):

I think it is C

OpenStudy (anonymous):

because mh is slightly bigger than kp

OpenStudy (anonymous):

I completely forgot to attach the picture lol

OpenStudy (anonymous):

yeah once I saw it I was able to back my decision 95%

OpenStudy (anonymous):

yeah once I saw it I was able to back up my decision by 95%

OpenStudy (anonymous):

LOL

OpenStudy (anonymous):

haha so why do you think it's C

OpenStudy (mathstudent55):



OpenStudy (anonymous):

I think it is C because 27 is slightly bigger than Mh which is 27.

OpenStudy (anonymous):

24 LOL

OpenStudy (mathstudent55):

To solve problems with similar triangles, it's good to see which sides are corresponding, and which sides we know and are being asked to find.

OpenStudy (anonymous):

LK Lol

OpenStudy (anonymous):

so i have to do a proportion

OpenStudy (anonymous):

yes I think So

OpenStudy (anonymous):

i got 48

OpenStudy (mathstudent55):

We want x. x corresponds to 18. We know 36, and 36 corresponds to 24. That allows us to write a proportion. \(\Large \dfrac{36}{24} = \dfrac{x}{18} \) Now we reduce the fraction on the left side, and we cross multiply.

OpenStudy (anonymous):

so it is 27

OpenStudy (anonymous):

i wish i could give medals to you both

OpenStudy (anonymous):

Yes it is 27

OpenStudy (mathstudent55):

\(\Large \dfrac{3}{2} = \dfrac{x}{18} \) \(\Large 2x = 3(18)\) \(\Large 2x = 54\) \(\Large x = 27\) Yes, you are correct.

OpenStudy (anonymous):

Don't give me the medal I just want to help

OpenStudy (anonymous):

i have more so stay tuned lol!!

OpenStudy (anonymous):

Ok just tag Me. :) LOL